OpenStudy (anonymous):

Average value of a function on an interval. Will post equstion and work in a comment

OpenStudy (anonymous):

\[\frac{ 1 }{ \pi - 0 } \int\limits\limits_{0}^{\pi} \sin (x) dx\] \[\int\limits_{0}^{\pi} \sin(x) dx = [-\cos] _{0}^{\pi}\] \[[-\cos \pi] - [-\cos 0] => [-(-1)] - [-1] = 1 + 1 = 2\] Average value = \[\frac{ 1 }{ \pi } (2) = \frac{ 2 }{ \pi }\] I'm asked to find the x value that corresponds to 2/pi sin(x) = 2/pi using the arcsin on the calculator I get ~ 0.690 But there is a second x value that corresponds to 2/pi, which is 2.451, how do I solve for the 2nd x value?

OpenStudy (anonymous):

Can I find the 2nd x value algebraically, using calculus, or using a graphing calculator?

OpenStudy (javalos01):

@rushwer

OpenStudy (javalos01):

@Rushwr

OpenStudy (irishboy123):

it's going to be symmetrical about \(\pi/2\)